# Estimating Spatially Resolved Radiation Fields Using Neural Networks

### v2

@@ Abstract
We present an in-depth analysis on how to build and train neural networks to estimate the spatial distribution of scattered radiation fields for radiation protection dosimetry in medical radiation fields, such as those found in interventional radiology and cardiology. We present three different synthetically generated datasets with increasing complexity for training, using a Monte-Carlo Simulation application based on Geant4. On those datasets, we evaluate convolutional and fully connected architectures of neural networks to demonstrate which design decisions work well for reconstructing the fluence and spectra distributions over the spatial domain of such radiation fields. All our datasets, as well as our training pipeline, are published as open source in separate repositories.
Code: https://github.com/Centrasis/radfield3d-nn
@@
Interventional Radiology, Monte-Carlo Simulation, FCNN, U-Net, Dataset, X-ray
@@
Monte-Carlo Simulation
@@
Dose-Structure Report
@@ 1 Introduction
In the context of IR, medical staff are exposed to inhomogeneous radiation fields due to their proximity to the patient, which complicates the accurate assessment of individual doses. The reliability of current personal dosimetry methods is contingent upon the uniform distribution of radiation fields. However, this assumption of uniformity does not hold in IR settings. To address this issue, previous research has proposed the use of computational dosimetry systems [1], which are designed to monitor the locations and postures of all individuals involved in an IR procedure. In order to assess doses in these exposure scenarios, such systems make use of currently available methods, that are mainly based on Monte-Carlo Simulations [2].
@@
To advance radiation protection, it is necessary to develop accelerated methods for the simulation of radiation transport. In this research work, we explore the surrogate model approach for MCS acceleration, whereby we learn spatially resolved radiation field distributions from a Monte-Carlo Simulation. Subsequently, the neural networks serve as efficient models for radiation-field reconstruction, providing a real-time capable radiation-transport simulation method. Furthermore, we have generated and made available a variety of datasets for the training of our models and analogous ones. The generation of our datasets was facilitated by the extension and utilization of the RadField3D [5] simulation software which itself is built upon the Geant4 [6] MCS framework to ensure physical correctness, as Geant4 is well established and validated for the targeted use-case.
@@ 1.1 Monte-Carlo Simulation in Radiation Protection
The technique of using MCS to calculate the radiation transport of complex measurements is already widely spread in various disciplines of dosimetry, from radiation protection to medical physics. Therefore, a set of well-established and tested general-purpose MCS toolkits for radiation transport already exists, such as Geant4 [6], MCNP [7], and EGSnrc [8]. As these toolkits are meant for general purpose applications ranging from astrophysics over nuclear reactor physics to dosimetry calculations, fine-tuning the MCS parameters for each specific use case is required. There are frameworks built upon these toolkits, such as GATE [9], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models based on the real tracking of personnel movements [10]. However, as highlighted by Rondon et al., current MCS capabilities are still far from real-time as they require, at best, several tens of seconds if not minutes to be performed. This is particularly challenging in IR, as the calculation of radiation-fluence distributions in the space around the patient during interventions is computationally demanding and can require hours of simulations. However, real-time simulation capabilities are becoming increasingly necessary for computational radiation protection dosimetry of staff, as proposed by prior research projects, such as the PODIUM project [1].
@@ 2 Materials and Methods
As there is a lack of publicly available, and in this context, usable datasets our first contribution is the generation and provision of three datasets for training spatially resolved dose rate predicting neural networks. For these three datasets, we implemented two distinct neural network architectures with additional variants for each, in order to analyze the influence of these architectures and specific hyperparameters on the quality of the predictions. For this purpose, we considered two basic architectures: an FCNN architecture inspired by Neural Radiance Fields [24] and a 3D-U-Net [25] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
@@ 2.1 Datasets
We generated and provided three synthetic datasets of radiation fields for different situations with increasing parameter spaces. The datasets were all generated using the RadField3D [5] dataset generator. The resulting datasets consist of one radiation field per fixed set of parameters. These fields are stored in the RadFiled3D file format and contain three channels. One for the photons of the direct X-ray beam, one for the resulting scattered fields and one containing the binary voxel occupation of the modeled geometry. Each of the first two channels contains the following information per voxel: The simulated volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$, the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ and the statistical error $\epsilon_{rel}$ as described in the research work about the MCS application RadField3D. Additionally, all input parameters, such as the X-ray tubes output spectra or the tubes position and direction, are stored per field.
@@
The first dataset we generated, contains $1250$ radiation fields. Each field has a resolution of 50×50×50 voxels with each voxel having an extent of $2\,\mathrm{cm}$ in every dimension resulting in a total field extent of $1\,\mathrm{m}$ per dimension. The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tubes output is fixed to a reference radiation quality of H-100 according to ISO 4037-1 [26] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$. The distance from the X-ray tube relative to the origin is $2.5\,\mathrm{m}$. At the origin of the field, the headless torso of a male Alderson RANDO phantom is placed to act as the scatter object for the radiation beam. This setting is derived from the validation measurements conducted for the RadField3D data generator [5], and, although simplified, it is considered realistic for modeling this first exposure scenario.
@@
The second dataset contains $2156$ radiation fields and extends the previous one by the radiation source spectra as a varying parameter. Here we used random, but realistic, spectra for modeling primary radiation in the context of IR. The range of the parameters was defined based on empirical feedback from medical physicist and radiologists, mainly gained during the PODIUM project. Therefore, we defined an energy range for the tube output energy $E_{\mathrm{tube}}\in[40,125]\,\mathrm{keV}$. Further, we added aluminum with a random thickness $t_{\mathrm{Al}}\in[2.5,7.5]\,\mathrm{mm}$ and, optionally, copper, with a random-thickness $t_{\mathrm{Cu}}\in[0.0,0.9]\,\mathrm{mm}$ for simulating realistic filtering material used to improve the X-ray beam quality. As the anode material, we set tungsten, which is a standard reference material in IR imaging devices, and used a random anode angle of $\angle_{\mathrm{Anode}}\in[8,12]\degree$. From that parameter space for the radiation source, we sampled sets of parameters and passed them to SpekPy [27] to generate reasonable X-ray spectra for the X-ray beams in this dataset.
@@
In the third and most complex dataset, dynamic X-ray tube distancing is introduced. A total of $3779$ distinct radiation fields were generated. For this dataset, we carried over the same base parameter space as described in the preceding datasets, and extended it by sampling the X-ray tube distance from the scenes isocenter as $d_{\mathrm{tube}}\in[35,75]\,\mathrm{cm}$. In order to make the dataset even more realistic in representing typical real-world scenarios, the X-ray beam was collimated to a rectangular shape, with fixed dimensions of $40\,\mathrm{cm}\times 30\,\mathrm{cm}$ in the plane passing through the origin.
@@
${DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}})}{\min(\Phi^{3}_{\mathrm{\gamma}})}\Bigr)$
@@
where ${DR}_{\mathrm{dB}}$ is used to describe the actual ranges needed for reconstruction and $G_{\mathrm{Gini}}$ expresses the deviation of the present fluence distribution from a uniform distribution. For the calculation of $G_{\mathrm{Gini}}$, the fluences are expected to be sorted in ascending order. A Gini coefficient of $G_{\mathrm{Gini}}=0$ indicates, that the dataset is uniformly distributed, while $G_{\mathrm{Gini}}=1$ indicates, that the dataset contains only one value or category.
@@
2.2 Data normalization
Figure 1: Distribution of relative voxel fluences inside datasets DS-01 and DS-02 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 0.0$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 10^{-2}$.
Figure 2: Distribution of relative voxel fluences inside datasets DS-01 and DS-02 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 10^{-1}$ on the X-axis, exhibiting a much weaker imbalance in that range.
Figure 3: Distribution of relative voxel fluences inside dataset DS-03 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 0.0$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 10^{-5}$.
Figure 4: Distribution of relative voxel fluences inside dataset DS-03 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 10^{-1}$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 20^{-1}$ and a small peak at $\Phi^{3}_{\mathrm{\gamma}}=1$.
In principle, radiation fields inside the dataset are already normalized, as the photon spectra are normalized to probability distributions $p(E_{\mathrm{\gamma}})$ and the fluences are normalized per single primary photon from the simulation. However, due to the numerical nature of the datasets, after loading a single radiation field from a dataset, we performed an extra normalization step for the data to limit the occurring gradients and avoid numerical instability, as well as introducing bias because of high total values. Thus, this extra normalization step effectively mitigates the severe class imbalance in all our dataset towards low fluence voxels. This has been visualized in the global fluence histograms in Fig. 4 for DS-01 and DS-02 and in Fig. 4 for DS-03. The histograms were generated using the MaxNorm function to transform the fluences to $[0,1]$. From both figures we can draw that the majority of voxels have a fluence lower than $5\,\%$ of the maximum fluence, which makes sense, as the major part of the radiation field is indeed scattered radiation. However, this imbalance could potentially lead to a mis-reconstruction of the primary beam. Especially for DS-03 from Fig. 4, we can see that the class imbalance is also present inside the primary beam component. Not as severely expressed, datasets DS-01 and DS-02 exhibit an unequal distribution of fluences as well, as visible from Fig. 4. Therefore, we examine the influence of different normalization functions regarding the reconstruction capabilities, to find the best fit to reconstruct the high and low fluence regions with sufficient results.
$\mathrm{MaxNorm}(x)$ $=\frac{x}{\max(\mathbf{X})},$ $\quad x\in\mathbf{X}\subset\mathbb{R}^{+}$
$\Rightarrow\mathrm{MaxNorm}:\mathbb{R}^{+}\rightarrow[0,1]$
$\mathrm{MaxLogNorm}_{\alpha}(x)$ $=\frac{\ln(1+\alpha\,x)}{\max_{y\in\mathbf{X}}(\ln(1+\alpha\,y))},$ $\quad x\in\mathbf{X}\subset\mathbb{R}^{+}$
$\Rightarrow\mathrm{MaxLogNorm}:\mathbb{R}^{+}\rightarrow[0,1]$
The idea of using alternative normalizations in the interval of $[0,1]$ is to create a more equal distribution of gradient strengths over the range of values to reconstruct. We hypothesize, that this is of importance for our use case, as the dosimetry is regarding relative instead of absolute errors per voxel. In general, it can be beneficial to elevate differences between small fluences and flatten the differences between higher fluences to give small deviations on small scales more impact. A classical solution for this issue is to use a transformation to the log-space, which is in general a reasonable choice for the following training of neural networks. Unfortunately, due to numerical instabilities, the steep ascent of the logarithm at very small values can become unstable under inversion and thus break the bijective property required for normalization. This instability can be observed for fluences of a magnitude lower than $10^{-8}$. As it can be seen from Tab. 1, for DS-03 fluences of about $10^{-8}$ play a role for the reconstruction as well. To mitigate this issue, we added a scaling factor $\alpha$ prior to the log-transformation. We tested two values for $\alpha$: $10^{0}$ and $10^{3}$. As an alternative to encounter the high dynamic range, we doubled the range of our previously linearly spread-out normalization from $\Phi^{3}_{\mathrm{\gamma}}\in[0,1]$ to $\Phi^{3}_{\mathrm{\gamma}}\in[-1,1]$ which additionally separates the low fluence voxels from the few high energy voxels by the sign. We distinguish between those two variants of the MaxNorm by indicating them as $\mathrm{MaxNorm}_{0}^{1}$ and $\mathrm{MaxNorm}_{-1}^{+1}$.
@@ 2.3 Architectures
In general, our models all estimate the relative topology of the spatial radiation field distribution conserving the relation of intensities between all voxels inside one specific radiation field. This is important, as the actual fluence scales between different settings of the beam would introduce a bias towards high X-ray tube currents and thus harm the training of the neural networks. On the downside, in a real-world application, this inherently requires our approach to run concurrently to at least one area dosemeter that reports dose rates at a known point in space. As these devices, especially those with real-time measurement capabilities, are known to exhibit a certain energy and thus spectrum dependence, one needs to correct them accordingly. This is specifically of importance in our targeted use case, as we aim to predict the scattered field around the patient, where the spectrum is not homogeneous and is difficult to predict. Thus, we decided to train all our networks not using the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ free-in-air or $H^{*}(10)$, but on the components of those measurands as provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$. In fact, the concurrent measurements of an area dosemeter are only needed to correct the fluence. Therefore, one can potentially use the estimated spectra to correct an area dosemeter measurement in turn.
@@
Our network begins by feeding the location vector through a encoder using frequency encoding using sine and cosine functions with different periodicities which were initially introduced by A. Vaswani et al.[31] for the transformer networks and reused for the original NeRF implementation. Even though, the direction vector can also be represented by a simple frequency encoding, we instead used a spherical harmonics based encoding as the cyclic nature of a rotation angle can be better represented by that encoding. This was demonstrated for NeRF-like networks by InstantNGP of T. Müller et al.[32]. For efficiency reasons, we used the spherical harmonics implementation as provided by the tiny-cuda-nn package [33] with version 2.0. For both encodings, we appended the raw input to the encoded vector. In contrast to the classical NeRF application, our network does not aim to learn the scatter objects 3D geometry with visual densities for volumetric rendering. Thus, we skipped the pre-processing path where the model should optimize the voxel location representation towards the local density independently from the direction. We want the network to learn the implicit representation of the fluence distribution, which is similar to the visual density, but in our case is ultimately dependent on the radiation direction. Thus, we extracted the fluence as well as the spectrum from the last layer. In order to build a robust backbone for solving the radiation field estimation under more complex parameter spaces, we did not just concatenate the feature vectors that describe the global beam conditions, such as the view direction in NeRF. Instead, we experimented with various feature fusing methods and finally compared the Feature-wise Linear Modulation (FiLM) [34] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [35] approach and as the baseline, with the simple concatenation of feature vectors as is done in classical NeRF networks. Ahead of fusing both feature vectors, the encoded X-ray tube direction vector was passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
The actual hyperparameters used for this and the extensions of this network, such as the frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusing layers are reported in the results section as we determined them using hyperparameter tuning. Through all layers of the main path of this network, except for the last ones, we use SiLU activation functions. Only in the case of the output activation did we use different activation functions. For the fluence, we used a different activation function depending on the range of the fluence normalization. For the case of a normalization in $[0,1]$, we used a sigmoid $\sigma$ activation function, while for a normalization in [1,1][-1,1], we used a gradient conserving clamping $\mathrm{clamp}_{\mathrm{GC}}$, with:
@@
${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}(x)=x+(\mathrm{clamp}_{min}^{max}(x)-x)$
For the gradient conserving clamping, the second term of the sum is executed detached from pytorchs autograd graph which results in a clamping that still has the original gradients attached. For the local spectra, we use a simple histogram normalization $\mathrm{norm}_{\mathrm{hist}}$ as the activation function, to guarantee normalized spectra for the losses.
@@
Our Beam2Scatter (B2S) U-Net is in general a CNN, that encodes the raw radiation field of the direct beam prior to the interaction with the patient geometry and from that estimates the scattered radiation field as it is induced by the patient. Therefore, its encoder-decoder architecture follows the 3D-U-Net description [25], which itself is just an adaption to the third dimension of the original U-Net, as it was initially used for biomedical image segmentation [36]. We adopted this structure and gave our network the fluence map of the direct beam as a one-channel input and let the network predict a $33$ dimensional output with the first $32$ channels for the spectra and the last channel for the fluence. Another addition is, that we introduced FiLM [34] layers after each up-sampling step in the decoder path, to modulate the weights of the decoder depending on the encoded input spectrum of the X-ray tube, which is encoded the same way as in SPERFNet. This way we can inject that information into the U-Net structure. The networks outputs had the prior introduced activation functions applied: ${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}$ and $\mathrm{norm}_{\mathrm{hist}}$. Thanks to the direct beam radiation distribution being the input of this network, it can be used directly with all three datasets. For DS-01, the spectra encoding and fusing by using FiLM layers was disabled.

### v3

# Learning-Based Estimation of Spatially Resolved Scatter Radiation Fields in Interventional Radiology

@@ Abstract
We present three variants of a lightweight, fully connected artificial neural network, suited for interactive estimation of three-dimensional, spatially resolved volumes of scattered radiation fields and a corresponding training pipeline for radiation protection dosimetry in medical radiation fields, such as those found in interventional radiology and cardiology. Accompanying, we present three different synthetically generated datasets with increasing complexity for training, generated using RadField3D, a Monte Carlo simulation application based on Geant4. As the primary scatter object, we employed the torso of a male Alderson RANDO phantom. On those datasets, we evaluate convolutional and fully connected architectures of neural networks to demonstrate which design decisions work well for reconstructing the fluence and spectra distributions over the spatial domain of such radiation fields. All our datasets, as well as our training pipeline, are published as open source in separate repositories.
To evaluate the presented neural networks, we define and assess several metrics. Across these measures, the model variants demonstrate good spatial agreement between predicted and ground-truth radiation fields, particularly within specific regions of interest within the radiation field. Of particular relevance for potential application in out-of-field dosimetry is the SMAPE of the scatter radiation field, which represents the most challenging metric and was consistently above $84\,\%$.
@@
Interventional Radiology, Monte Carlo Simulation, FCNN, U-Net, Dataset, X-ray
@@
Monte Carlo Simulation
@@
Dose Structured Report
@@ 1 Introduction
In the context of IR, medical staff are exposed to inhomogeneous radiation fields due to their proximity to the patient, which complicates the accurate assessment of individual doses. The reliability of current personal dosimetry methods is contingent upon the uniform distribution of radiation fields. However, this assumption of uniformity does not hold in IR settings. To address this issue, previous research has proposed the use of computational dosimetry systems [21], which are designed to monitor the locations and postures of all individuals involved in an IR procedure. In order to assess doses in these exposure scenarios, such systems make use of currently available methods, that are mainly based on Monte Carlo Simulations [6].
@@
To advance radiation protection, it is necessary to develop accelerated methods for the simulation of radiation transport. In this research work, we explore the surrogate model approach for MCS acceleration, whereby we learn spatially resolved radiation field distributions from a Monte Carlo Simulation. Subsequently, the neural networks serve as efficient models for radiation-field reconstruction, providing a real-time capable radiation-transport simulation method. Furthermore, we have generated and made available a variety of datasets for the training of our models and analogous ones. The generation of our datasets was facilitated by the extension and utilization of the RadField3D [15] simulation software which itself is built upon the Geant4 [2] MCS framework to ensure physical correctness, as Geant4 is well established and validated for the targeted use-case.
@@ 1.1 Monte-Carlo Simulation in Radiation Protection
The technique of using MCS to calculate the radiation transport of complex measurements is already widely spread in various disciplines of dosimetry, from radiation protection to medical physics. Therefore, a set of well-established and tested general-purpose MCS toolkits for radiation transport already exists, such as Geant4 [2], MCNP [40], and EGSnrc [22]. As these toolkits are meant for general purpose applications ranging from astrophysics over nuclear reactor physics to dosimetry calculations, fine-tuning the MCS parameters for each specific use case is required. There are frameworks built upon these toolkits, such as GATE [13], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models based on the real tracking of personnel movements as demonstrated by Rondón et al. [30]. However, as highlighted by Rondón et al., current MCS capabilities are still far from real-time as they require, at best, several tens of seconds if not minutes to be performed. This is particularly challenging in IR, as the calculation of radiation-fluence distributions in the space around the patient during interventions is computationally demanding and can require hours of simulations. However, real-time simulation capabilities are becoming increasingly necessary for computational radiation protection dosimetry of staff, as proposed by prior research projects, such as the PODIUM project [21].
@@ 1.3 Hybrid Methods and Surrogate Models for Radiation Transport MCS
This work aims to contribute to the corpus of existing surrogate model approaches by focusing on predicting dynamic radiation fields using lightweight FCNNs. This enables the prediction of partial radiation fields within the region of interest, thereby supporting real-time capable radiation field estimation and, in theory, also allowing for higher resolutions of radiation fields. In contrast to prior approaches, our approach additionally provides the local energy spectra, making it possible to use those networks for correcting the energy dependence of real world dosemeters. In the following, we present FCNNs inspired by the architecture of the networks used for Neural Radiance Fields [18]. As these networks model geometry implicitly, and our datasets do not incorporate variations in patient geometry, a direct comparison with the work of Arias et al. [35] is neither meaningful nor feasible, given that their approach explicitly encodes patient geometry.
@@ 2 Materials and Methods
As there is a lack of publicly available, and in this context, usable datasets our first contribution is the generation and provision of three datasets for training spatially resolved dose rate predicting neural networks. For these three datasets, we implemented two distinct neural network architectures with additional variants for each, in order to analyze the influence of these architectures and specific hyperparameters on the quality of the predictions. For this purpose, we considered two basic architectures: an FCNN architecture inspired by NeRFs [18] and a 3D-U-Net [7] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
@@ 2.1 Datasets
We generated and provided three synthetic datasets of radiation fields for different situations with increasing parameter spaces. The datasets were all generated using the RadField3D [15] dataset generator. Per field, the generation took about $20\,\mathrm{min}$ on our AMD Epyc 9474F 48-core CPU. The resulting datasets consist of one radiation field per fixed set of parameters. These fields are stored in the RadFiled3D file format and contain three channels. One for the photons of the direct X-ray beam, one for the resulting scattered fields and one containing the binary voxel occupation of the modeled geometry. Each of the first two channels contains the following information per voxel: The simulated volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ (fluence per cubic volume of a voxel), the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ and the statistical error $\epsilon_{rel}$ as described in the research work about the MCS application RadField3D. Additionally, all input parameters, such as the X-ray tube output spectra or the tubes position and direction, are stored per field.
@@
The first dataset we generated, contains $1250$ radiation fields. Each field has a resolution of 50×50×50 voxels with each voxel having an extent of $2\,\mathrm{cm}$ in every dimension resulting in a total field extent of $1\,\mathrm{m}$ per dimension. The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tube output is fixed to a reference radiation quality of H-100 according to ISO 4037-1 [9] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$. The distance from the X-ray tube relative to the origin is $2.5\,\mathrm{m}$. At the origin of the field, the headless torso of a male Alderson RANDO phantom [27] is placed to act as the scatter object for the radiation beam. This setting is derived from the validation measurements conducted for the RadField3D data generator [15], and, although simplified, it is considered realistic for modeling this first exposure scenario.
@@
The second dataset contains $2156$ radiation fields and extends the previous one by the radiation source spectra as a varying parameter. Here we used random, but realistic, spectra for modeling primary radiation in the context of IR. The range of the parameters was defined based on empirical feedback from medical physicist and radiologists, mainly gained during the PODIUM project. Therefore, we defined an energy range for the tube output peak energy $E_{\mathrm{tube}}\in[40,125]\,\mathrm{keV}$. Further, we added aluminum with a random thickness $t_{\mathrm{Al}}\in[2.5,7.5]\,\mathrm{mm}$ and, optionally, copper, with a random-thickness $t_{\mathrm{Cu}}\in[0.0,0.9]\,\mathrm{mm}$ for simulating realistic filtering material used to improve the X-ray beam quality. As the anode material, we set tungsten, which is a standard reference material in IR imaging devices, and used a random anode angle of $\angle_{\mathrm{Anode}}\in[8,12]\degree$. From that parameter space for the radiation source, we sampled sets of parameters and passed them to SpekPy [25] to generate reasonable X-ray spectra for the X-ray beams in this dataset.
@@
In the third and most complex dataset, dynamic X-ray tube distancing is introduced. A total of $3779$ distinct radiation fields were generated. For this dataset, we mostly carried over the same base parameter space as described in the preceding datasets, but in contrast to the previous datasets, we sampled the X-ray tube distance from the scenes isocenter as $d_{\mathrm{tube}}\in[35,75]\,\mathrm{cm}$. In order to make the dataset even more realistic in representing typical real-world scenarios, the X-ray beam was collimated to a rectangular shape, with fixed dimensions of $40\,\mathrm{cm}\times 30\,\mathrm{cm}$ in the plane passing through the origin and thus intersecting the phantom.
Table 1: Listing of the statistical errors, with respect to the fluence, from the scattered radiation fields separated by dataset. $\mathrm{mean}(error)$ refers to the mean over all voxels in the dataset, while $\mathrm{median}(error)$ refers to the mean over the median statistical error of each file inside the dataset.
Dataset $\mathrm{mean}(error)$ $\mathrm{median}(error)$
DS-01 $9.7\,\%$ $0.03\,\%$
DS-02 $9.9\,\%$ $0.03\,\%$
DS-03 $9.9\,\%$ $0.03\,\%$
RadField3D uses the variation of the relative voxel fluences during the simulation process to measure the statistical error per voxel. Those statistical errors from the scattered radiation field layer are evaluated continuously to determine the current overall statistical error. Once $90\,\%$ of those voxels exhibit a statistical error equal to or less than $10\,\%$, the simulation is aborted. This behavior is reflected by Tab. 1 where we show the mean and median statistical errors present in our previously described datasets, as calculated by RadField3D.
An example of the spatial distribution of relative air kerma rates in air $\dot{K}_{\mathrm{air}}$ can bee seen in Fig. 1. The two projections were generated from a random field taken from DS-03 by excluding all doserates below $0.5\,\%$ of the maximum doserate for improved visual understanding.
Figure 1: Two perspective projections of a random radiation field from DS-03: One projection along the X-axis (left) and one along Y-axis (right). The spatially resolved relative doserate $\dot{K}_{\mathrm{air}}$ (air kerma in air) is color coded from transparent ($0.5\,\%$) over blue ($0.5\,\%$) to yellow ($100\,\%$). The used scatter geometry (Male Alderson RANDO phantom torso) is colored in red.
@@
${DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}}\penalty 10000\ )}{\min(\Phi^{3}_{\mathrm{\gamma}}\penalty 10000\ )}\Bigr)$
@@
where ${DR}_{\mathrm{dB}}$ is used to describe the actual ranges needed for reconstruction and $G_{\mathrm{Gini}}$ expresses the deviation of the present fluence distribution from a uniform distribution. For the calculation of $G_{\mathrm{Gini}}$, the fluences are expected to be sorted in ascending order , where $\mu$ denotes the mean fluence and $n$ the total count of voxels across the dataset. Finally, a Gini coefficient of $G_{\mathrm{Gini}}=0$ indicates that the dataset is uniformly distributed, while $G_{\mathrm{Gini}}=1$ indicates that all fluences of the dataset are accumulated in one single bin of the fluence histogram.
@@
Overall, for all three datasets, the dataset generator RadField3D ensures, that $90\,\%$ of all voxels do not exceed a statistical error of $10\,\%$ with respect to fluence.
Data normalization:
All datasets used are provided normalized by RadField3D. This applies both to the voxelwise photon energy spectra, which are probability distributions with unit integral, and to the voxelwise fluences, which are normalized with respect to the number primary particles of the simulation. This effectively scales all fluence values relative to the maximum, ensuring that all fluences fall within the $[0..1]$ interval.
@@ 2.2 Architectures
In general, our models all estimate the relative topology of the spatial radiation field distribution conserving the relation of intensities between all voxels inside one specific radiation field. This is important, as the actual fluence scales between different settings of the beam would introduce a bias towards high X-ray tube currents and thus harm the training of the neural networks. On the downside, in a real-world application, this inherently requires our approach to run concurrently to at least one area dosemeter that reports dose rates at a known point in space. As these devices, especially those with real-time measurement capabilities, are known to exhibit a certain energy and thus spectrum dependence, one needs to correct them accordingly. This is specifically of importance in our targeted use case, as we aim to predict the scattered field around the patient, where the spectrum is not homogeneous and is difficult to predict. Thus, we decided to train all our networks not using the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ free-in-air or ambient dose equivalent $H^{*}(10)$, but on the components of those measurands as provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$. In fact, the concurrent measurements of an area dosemeter are only needed to correct the fluence. Therefore, one can potentially use the estimated spectra to correct an area dosemeter measurement for its energy and angle dependence in turn.
@@
Our network begins by feeding the location vector through an encoder using frequency encoding composed of sine and cosine functions with different periodicities which were initially introduced by A. Vaswani et al.[34] for the transformer networks and reused for the original NeRF implementation. Even though, the direction vector can also be represented by a simple frequency encoding, we instead used a spherical harmonics based encoding as the cyclic nature of a rotation angle can be better represented by that encoding. This was demonstrated for NeRF-like networks by InstantNGP of T. Müller et al.[19]. For efficiency reasons, we used the spherical harmonics implementation as provided by the tiny-cuda-nn package [20] with version 2.0. For both encodings, we appended the raw input to the encoded vector. In contrast to the classical NeRF application, our network does not aim to learn the scatter objects 3D geometry with visual densities for volumetric rendering. Thus, we skipped the pre-processing path where the model should optimize the voxel location representation towards the local density independently from the direction. We want the network to learn the implicit representation of the fluence distribution, which is similar to the visual density, but in our case is ultimately dependent on the radiation direction. Thus, we extracted the fluence as well as the spectrum from the last layer. In order to build a robust backbone for solving the radiation field estimation under more complex parameter spaces, we did not just concatenate the feature vectors that describe the global beam conditions, such as the view direction in NeRF. Instead, we experimented with various feature fusion methods and finally compared the Feature-wise Linear Modulation (FiLM) [24] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [4] approach and as the baseline, with the simple concatenation of feature vectors as is done in classical NeRF networks. Ahead of fusion both feature vectors, the encoded X-ray tube direction vector was passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
The actual hyperparameters used for this and the extensions of this network, such as the frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusion layers are reported in the results section as we determined them using hyperparameter tuning. Through all layers of the main path of this network, except for the last ones, we use Sigmoid Linear Unit (SiLU) activation functions. For the output layer, we used two different activation functions. For the fluence, we used the sigmoid activation, to keep the output between 0 and 1:
@@
For the local spectra, we use a simple histogram normalization $\mathrm{norm}_{\mathrm{hist}}$ as the activation function, to guarantee normalized spectra for the losses.
@@
Our Beam2Scatter (B2S) U-Net is in general a CNN, that encodes the raw radiation field of the direct beam prior to the interaction with the patient geometry and from that estimates the scattered radiation field as it is induced by the patient. Therefore, its encoder-decoder architecture follows the 3D-U-Net description [7], which itself is just an adaption to the third dimension of the original U-Net, as it was initially used for biomedical image segmentation [29]. We adopted this structure and gave our network the fluence map of the direct beam as a one-channel input and let the network predict a $33$ dimensional output with the first $32$ channels for the spectra and the last channel for the fluence. Another addition is, that we introduced FiLM [24] layers after each up-sampling step in the decoder path, to modulate the weights of the decoder depending on the encoded input spectrum of the X-ray tube, which is encoded the same way as in SPERFNet. This way we can inject that information into the U-Net structure. The networks outputs had the prior introduced activation functions applied: sigmoid and $\mathrm{norm}_{\mathrm{hist}}$. Thanks to the direct beam radiation distribution being the input of this network, it can be used directly with all three datasets. For DS-01, the spectra encoding and fusion by using FiLM layers was disabled.